# Two-phase flows through porous media described by a Cahn–Hilliard–Brinkman model with dynamic boundary conditions

### v1

@@ 2.5.2   Definition of weak solutions for singular potentials
Then, $({\boldsymbol{v}},\varphi,\xi,\mu,\psi,\xi_{\Gamma},\theta)$ is called a weak solution of the Cahn–Hilliard–Brinkman system (1) if the following conditions are fulfilled:
@@
    $(\varphi,\psi)$ $\in H^{1}(0,T;{\cal V}_{K}^{*})\cap C^{0}([0,T];{{\cal H}})\cap L^{\infty}(0,T;{\cal V}_{K}),\quad$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 2.5.3   Existence of a weak solution
Let $K\geq 0$ be arbitrary. Suppose that (A1)(A3) and (S1)(S2) are fulfilled. Let $(\varphi_{0},\psi_{0})\in{\cal V}_{K}$ denote any initial data satisfying (2.8). In the case $K=0$, let the domain $\Omega$ be of class $C^{2}$. Then, the Cahn–Hilliard–Brinkman system (1) admits at least one weak solution $({\boldsymbol{v}},\varphi,\xi,\mu,\psi,\xi_{\Gamma},\theta)$ in the sense of Definition 2.8. In all cases, if the domain $\Omega$ is at least of class $C^{2}$, it holds that
@@
$\theta$ $=-\Delta_{\Gamma}\psi+\xi_{\Gamma}+\pi_{\Gamma}(\psi)+\partial_{\bf n}\varphi\quad$ on $\text{on $\Sigma$},$ (2.31)
@@ 3.1   Existence of weak solutions in the case $K>0$
We intend to construct a weak solution to system (1) by discretizing the weak formulation ((ii)) by means of a Faedo–Galerkin scheme. In this proof, the letter $C$ will denote generic positive constants that may depend on $K$, $\Omega$, $T$, the initial data and the constants introduced in (A1)(A3), and may change their value from line to line. Recall that, as $K$ is assumed to be positive here, we have $\sigma(K)=\frac{1}{K}$.
@@ 3.1.1   Construction of local-in-time approximate solutions
are dense. To construct a sequence of approximate solutions, we now make the ansatz
[…]
@@
Let now $k\in{\mathbb{N}}$ and $t\in[0,T]$ be arbitrary. We consider the bilinear form
$\mathcal{B}_{k,t}:{\bf V}_{\sigma,{\bf n}}\times{\bf V}_{\sigma,{\bf n}}\to{\mathbb{R}},$
$({\boldsymbol{v}},{\boldsymbol{w}})\mapsto 2\int_{\Omega}\nu\big{(}\varphi_{k}(t)\big{)}\,\mathbf{D}{\boldsymbol{v}}:\mathbf{D}{\boldsymbol{w}}+\int_{\Omega}\lambda\big{(}\varphi_{k}(t)\big{)}\,{\boldsymbol{v}}\cdot{\boldsymbol{w}}+\int_{\Gamma}\gamma\big{(}\psi_{k}(t)\big{)}\,{\boldsymbol{v}}\cdot{\boldsymbol{w}},$
which is related to the weak formulation of the Brinkman equation with Navier-slip boundary condition. It is obvious that $\mathcal{B}_{k,t}$ is symmetric, and, in view of (A3), it is easy to see that $\mathcal{B}_{k,t}$ is continuous. We further recall that every ${\boldsymbol{v}}\in{\bf V}_{\sigma,{\bf n}}$ satisfies $\mathop{\rm div}\nolimits({\boldsymbol{v}})=0$ a.e. in $\Omega$, and ${\boldsymbol{v}}\cdot{\bf n}=0$ a.e. on $\Gamma$. By means of (A3), integration by parts and the Poincaré inequality presented in Lemma 2.1, we deduce
$\mathcal{B}_{k,t}({\boldsymbol{v}},{\boldsymbol{v}})$ $\geq 2\nu_{1}\int_{\Omega}\,\mathbf{D}{\boldsymbol{v}}:\mathbf{D}{\boldsymbol{v}}+\gamma_{1}\int_{\Gamma}|{\boldsymbol{v}}|^{2}$
$=\nu_{1}\int_{\Omega}|\nabla{\boldsymbol{v}}|^{2}+\gamma_{1}\int_{\Gamma}|{\boldsymbol{v}}|^{2}\geq C\|{\boldsymbol{v}}\|_{{\bf{V}}}^{2}$
@@
$\mathcal{B}_{k,t}\big{(}{\boldsymbol{v}}_{k}(t),{\boldsymbol{w}}\big{)}=-\int_{\Omega}\varphi_{k}(t)\nabla\mu_{k}(t)\cdot{\boldsymbol{w}}-\int_{\Gamma}\psi_{k}(t)\nabla_{\Gamma}\theta_{k}(t)\cdot{\boldsymbol{w}}$
@@
Choosing $\zeta=u_{j}$ in (3.4b) and $\zeta_{\Gamma}=v_{j}$ in (3.4c) for $j=1,...,k$, we infer that $(\mathbf{a}^{k},\mathbf{b}^{k})^{\top}$ is determined by a system of $2k$ nonlinear ordinary differential equations subject to the initial conditions
@@
for all $i\in\{1,...,k\}$. In particular, since the functions $M_{\Omega}$ and $M_{\Gamma}$ are continuous and ${\boldsymbol{v}}_{k}$ depends continuously on the coefficients $\mathbf{a}^{k}$, $\mathbf{b}^{k}$, $\mathbf{c}^{k}$ and $\mathbf{d}^{k}$, the same holds for the right-hand side of this ODE system. Moreover, choosing $(\eta,\eta_{\Gamma})=(u_{j},0)$ and $(\eta,\eta_{\Gamma})=(0,v_{j})$ for $j=1,...,k$ in (3.1.1), respectively, we find that the coefficients $\mathbf{c}^{k}$ and $\mathbf{d}^{k}$ are explicitly given by $2k$ algebraic equations whose right-hand side depends continuously on $\mathbf{a}^{k}$ and $\mathbf{b}^{k}$. This allows us to replace $\mathbf{c}^{k}$ and $\mathbf{d}^{k}$ in the right-hand side of the aforementioned ODE system to obtain a closed ODE system for the vector-valued function $(\mathbf{a}^{k},\mathbf{b}^{k})^{\top}$ whose right-hand side depends continuously on $(\mathbf{a}^{k},\mathbf{b}^{k})^{\top}$. Consequently, the Cauchy–Peano theorem implies the existence of at least one local-in-time solution
@@
to the corresponding initial value problem. Here, we take $T_{k}^{*}>0$ as large as possible meaning that $[0,T_{k}^{*})\cap[0,T]$ is the right-maximal time interval of this solution. We can now reconstruct
@@
${\boldsymbol{v}}_{k}\in C^{1}([0,T_{k}^{*});{\bf V}_{\sigma,{\bf n}})\quad\text{and}\quad(\varphi_{k},\psi_{k}),(\mu_{k},\theta_{k})\in C^{1}([0,T_{k}^{*});{\cal V}),$
which fulfills the discretized weak formulation (3.1.1) on the time interval $[0,T_{k}^{*})$.
@@ 3.1.2   Uniform estimates
First estimate. We first test (3.4a) by ${\boldsymbol{v}}_{k}$, (3.4b) by $\mu_{k}$, (3.4c) by $\theta_{k}$, and (3.1.1) by $-(\partial_{t}\varphi_{k},\partial_{t}\psi_{k})$. We then add these equations and integrate the resulting equation with respect to time from 00 to an arbitrary $t\in[0,T_{k}]$. We obtain
@@
$\left\|\mathbf{D}{\boldsymbol{v}}_{k}\right\|_{L^{2}(0,T_{k};{\mathbb{H}})}+\left\|{\boldsymbol{v}}_{k}\right\|_{L^{2}(0,T_{k};\mathbf{H}_{\Gamma})}+\left\|\nabla\varphi_{k}\right\|_{L^{\infty}(0,T_{k};{\bf{H}})}+\left\|\nabla_{\Gamma}\psi_{k}\right\|_{L^{\infty}(0,T_{k};{\bf{H}}_{\Gamma})}$
@@
Invoking the Poincaré inequality presented in Lemma 2.1, we directly infer
@@
Hence, in view of (3.7) and (3.1.2), we use the Poincaré–Wirtinger inequality to conclude
@@
$\quad\leq\left(\left\|\varphi_{k}\right\|_{L^{\infty}(0,T_{k};L^{4}(\Omega))}\left\|{\boldsymbol{v}}_{k}\right\|_{L^{2}(0,T_{k};(L^{4}(\Omega))^{d})}+M_{2}\left\|\nabla\mu_{k}\right\|_{L^{2}(0,T_{k};{\bf H})}\right)\left\|{\overline{\zeta}}\right\|_{L^{2}(0,T_{k};V)}$
@@
Third estimate. Next, we want to derive uniform bounds on $\mu_{k}$ in $L^{4}(0,T_{k};H)\cap L^{2}(0,T_{k};V)$ and on $\theta_{k}$ in $L^{4}(0,T_{k};{H}_{\Gamma})\cap L^{2}(0,T_{k};{V}_{\Gamma})$. Therefore, we choose arbitrary functions $\eta\in L^{1}(0,T_{k};V)$ and $\eta_{\Gamma}\in L^{1}(0,T_{k};{V}_{\Gamma})$ and we set ${\overline{\eta}}:=\mathbb{P}_{V_{k}}(\eta)$ and ${\overline{\eta_{\Gamma}}}:=\mathbb{P}_{V_{\Gamma,k}}(\eta_{\Gamma})$. Testing (3.1.1) by $({\overline{\eta}},{\overline{{\eta_{\Gamma}}}})$, recalling the growth conditions from (R1) as well as the uniform bounds (3.1.2) and (3.10), we use Hölder’s inequality and Sobolev’s embedding theorem to derive the estimate
$\int_{0}^{T_{k}}\big{|}\mathopen{\langle}(\mu_{k},\theta_{k}),(\eta,{\eta_{\Gamma}})\mathclose{\rangle}_{{\cal V}}\big{|}=\int_{0}^{T_{k}}\big{|}\mathopen{\langle}(\mu_{k},\theta_{k}),({\overline{\eta}},{\overline{{\eta_{\Gamma}}}})\mathclose{\rangle}_{{\cal V}}\big{|}$
$\quad\leq\int_{0}^{T_{k}}\Big{[}\left\|\nabla\varphi_{k}\right\|_{{\bf H}}\left\|\nabla{\overline{\eta}}\right\|_{{\bf H}}+\left\|F^{\prime}(\varphi_{k})\right\|_{L^{\frac{6}{5}}}(\Omega)\left\|{\overline{\eta}}\right\|_{L^{6}(\Omega)}+\left\|\nabla_{\Gamma}\psi_{k}\right\|_{{\bf H}_{\Gamma}}\left\|\nabla_{\Gamma}{\overline{{\eta_{\Gamma}}}}\right\|_{{\bf H}_{\Gamma}}$
@@
in $[0,T_{k}]$. Taking the supremum over all $(\eta,{\eta_{\Gamma}})\in L^{1}(0,T_{k};{\cal V})$ with $\left\|(\eta,{\eta_{\Gamma}})\right\|_{L^{1}(0,T_{k};{\cal V})}\leq 1$, and using (3.11), we infer

### v2

@@ 2.5.2 Definition of weak solutions for singular potentials
Then, $({\boldsymbol{v}},\varphi,\xi,\mu,\psi,\xi_{\Gamma},\theta)$ is called a weak solution of the Cahn–Hilliard–Brinkman system (1.1) if the following conditions are fulfilled:
@@
    $(\varphi,\psi)$ $\in H^{1}(0,T;{\cal V}^{*})\cap C^{0}([0,T];{{\cal H}})\cap L^{\infty}(0,T;{\cal V}_{K}),\quad$
@@
    $\begin{aligned} &2\int_{\Omega}\nu(\varphi)\mathbf{D}{\boldsymbol{v}}:\mathbf{D}{\boldsymbol{w}}+\int_{\Omega}\lambda(\varphi){\boldsymbol{v}}\cdot{\boldsymbol{w}}+\int_{\Gamma}\gamma(\psi){\boldsymbol{v}}\cdot{\boldsymbol{w}}\\ &\quad=-\int_{\Omega}\varphi\nabla\mu\cdot{\boldsymbol{w}}-\int_{\Gamma}\psi\nabla_{\Gamma}\theta\cdot{\boldsymbol{w}},\end{aligned}$ (2.32a)
@@
    $\begin{aligned} \int_{\Omega}\mu\eta+\int_{\Gamma}\theta{\eta_{\Gamma}}&=\int_{\Omega}\nabla\varphi\cdot\nabla\eta+\int_{\Omega}\xi\eta+\int_{\Omega}\pi(\varphi)\eta+\int_{\Gamma}\nabla_{\Gamma}\psi\cdot\nabla_{\Gamma}{\eta_{\Gamma}}\\ &\quad+\int_{\Gamma}\xi_{\Gamma}{\eta_{\Gamma}}+\int_{\Gamma}\pi_{\Gamma}(\psi){\eta_{\Gamma}}+\sigma(K)\int_{\Gamma}(\psi-\varphi)({\eta_{\Gamma}}-\eta)\end{aligned}$ (2.32d)
@@
    $\xi\in\beta(\varphi)\quad\hbox{a.e.\ in~$Q$},\quad\xi_{\Gamma}\in\beta_{\Gamma}(\psi)\quad\hbox{a.e.\ on~$\Sigma$}.$
@@
    $\varphi(0)=\varphi_{0}\quad\hbox{a.e.\ in~$\Omega$},\quad\psi(0)=\psi_{0}\quad\hbox{a.e.\ on~$\Gamma$}.$
@@
Notice that, if $\Omega$ is connected, the pressure $p$ can be reconstructed from (2.32a) by proceeding as in Remark 2.5.
@@ 2.5.3 Existence of a weak solution
Let $K\geq 0$ be arbitrary. Suppose that (A1)(A3) and (S1)(S2) are fulfilled. Let $(\varphi_{0},\psi_{0})\in{\cal V}_{K}$ denote any initial data satisfying (2.31). In the case $K=0$, let the domain $\Omega$ be of class $C^{2}$. Then, the Cahn–Hilliard–Brinkman system (1.1) admits at least one weak solution $({\boldsymbol{v}},\varphi,\xi,\mu,\psi,\xi_{\Gamma},\theta)$ in the sense of Definition 2.9. In all cases, if the domain $\Omega$ is at least of class $C^{2}$, it holds that
@@
$\theta$ $=-\Delta_{\Gamma}\psi+\xi_{\Gamma}+\pi_{\Gamma}(\psi)+\partial_{\bf n}\varphi\quad$ $\text{on $\Sigma$},$ (2.35)
@@ 3.1  Existence of weak solutions in the case $K>0$
We intend to construct a weak solution to system (1.1) by discretizing the weak formulation (2.16) by means of a Faedo–Galerkin scheme. In this proof, the letter $C$ will denote generic positive constants that may depend on $K$, $\Omega$, $T$, the initial data and the constants introduced in (A1)(A3), and may change their value from line to line. Recall that, as $K$ is assumed to be positive here, we have $\sigma(K)=\frac{1}{K}$.
@@ 3.1.1 Construction of local-in-time approximate solutions
are dense. In order to construct a sequence of approximate solutions, we use a semi-Galerkin approach. This means that only the quantities $\varphi$, $\psi$, $\mu$ and $\theta$ are approximated by a Galerkin-scheme, and the approximate velocity field is obtained by directly solving the corresponding Brinkman subsystem. This approach has already been employed in [35] for a Cahn–Hilliard–Brinkman model without dynamic boundary conditions and a no-friction boundary condition for the velocity equation. Compared to the present paper, some of the steps are carried out in [35] in more detail, so we recommend it as a reference work.
To construct a sequence of approximate solutions, we now make the ansatz
$\begin{aligned} \varphi_{k}(x,t)&=\sum_{i=1}^{k}a_{i}^{k}(t)u_{i}(x),\quad&&\psi_{k}(x,t)=\sum_{i=1}^{k}b_{i}^{k}(t)v_{i}(x),\\ \mu_{k}(x,t)&=\sum_{i=1}^{k}c_{i}^{k}(t)u_{i}(x),\quad&&\theta_{k}(x,t)=\sum_{i=1}^{k}d_{i}^{k}(t)v_{i}(x)\end{aligned}$ (3.3)
@@
Let now $k\in{\mathbb{N}}$ and $t\in[0,T]$ be arbitrary. For any choice of $\mathbf{a}^{k}$, $\mathbf{b}^{k}$, $\mathbf{c}^{k}$ and $\mathbf{d}^{k}$, and $(\varphi_{k},\psi_{k},\mu_{k},\theta_{k})$ as defined in (3.3), we consider the bilinear form
$\begin{aligned} &\mathcal{B}_{k,t}:{\bf V}_{\sigma,{\bf n}}\times{\bf V}_{\sigma,{\bf n}}\to{\mathbb{R}},\\ &({\boldsymbol{v}},{\boldsymbol{w}})\mapsto 2\int_{\Omega}\nu\big{(}\varphi_{k}(t)\big{)}\,\mathbf{D}{\boldsymbol{v}}:\mathbf{D}{\boldsymbol{w}}+\int_{\Omega}\lambda\big{(}\varphi_{k}(t)\big{)}\,{\boldsymbol{v}}\cdot{\boldsymbol{w}}+\int_{\Gamma}\gamma\big{(}\psi_{k}(t)\big{)}\,{\boldsymbol{v}}\cdot{\boldsymbol{w}},\end{aligned}$ (3.4)
which is related to the weak formulation of the Brinkman equation with Navier-slip boundary condition. It is obvious that $\mathcal{B}_{k,t}$ is symmetric, and, in view of (A3), it is easy to see that $\mathcal{B}_{k,t}$ is continuous. We further recall that every ${\boldsymbol{v}}\in{\bf V}_{\sigma,{\bf n}}$ satisfies $\mathop{\rm div}\nolimits({\boldsymbol{v}})=0$ a.e. in $\Omega$, and ${\boldsymbol{v}}\cdot{\bf n}=0$ a.e. on $\Gamma$. Now, if (A3)(A3.1) holds, we use Korn’s inequality (Lemma 2.1(a)) to deduce
$\mathcal{B}_{k,t}({\boldsymbol{v}},{\boldsymbol{v}})\geq 2\nu_{1}\int_{\Omega}\,\mathbf{D}{\boldsymbol{v}}:\mathbf{D}{\boldsymbol{v}}+\gamma_{1}\int_{\Gamma}|{\boldsymbol{v}}|^{2}\geq\min\{2\nu_{1},\gamma_{1}\}\big{(}\left\|\mathbf{D}{\boldsymbol{v}}\right\|_{{\mathbb{H}}}^{2}+\|{\boldsymbol{v}}\|_{\mathbf{H}_{\Gamma}}^{2}\big{)}\geq C\left\|{\boldsymbol{v}}\right\|_{{\bf{V}}}^{2}$ (3.5)
for all ${\boldsymbol{v}}\in{\bf V}_{\sigma,{\bf n}}$. On the other hand, if (A3)(A3.2) holds, we use Korn’s inequality (Lemma 2.1(b)) to conclude
$\mathcal{B}_{k,t}({\boldsymbol{v}},{\boldsymbol{v}})\geq 2\nu_{1}\int_{\Omega}\,\mathbf{D}{\boldsymbol{v}}:\mathbf{D}{\boldsymbol{v}}\geq C\left\|{\boldsymbol{v}}\right\|_{{\bf{V}}}^{2}$ (3.6)
@@
$\mathcal{B}_{k,t}\big{(}{\boldsymbol{v}}_{k}(t),{\boldsymbol{w}}\big{)}=-\int_{\Omega}\varphi_{k}(t)\nabla\mu_{k}(t)\cdot{\boldsymbol{w}}-\int_{\Gamma}\psi_{k}(t)\nabla_{\Gamma}\theta_{k}(t)\cdot{\boldsymbol{w}}$ (3.7)
@@
Choosing $\zeta=u_{j}$ in (3.8b) and $\zeta_{\Gamma}=v_{j}$ in (3.8c) for j=1,,kj=1,...,k, we infer that $(\mathbf{a}^{k},\mathbf{b}^{k})^{\top}$ is determined by a system of $2k$ nonlinear ordinary differential equations subject to the initial conditions
@@
for all $i\in\{1,...,k\}$. In particular, since the functions $M_{\Omega}$ and $M_{\Gamma}$ are continuous and ${\boldsymbol{v}}_{k}$ depends continuously on the coefficients $\mathbf{a}^{k}$, $\mathbf{b}^{k}$, $\mathbf{c}^{k}$ and $\mathbf{d}^{k}$, the same holds for the right-hand side of this ODE system. Moreover, choosing $(\eta,\eta_{\Gamma})=(u_{j},0)$ and $(\eta,\eta_{\Gamma})=(0,v_{j})$ for j=1,,kj=1,...,k in (3.8), respectively, we find that the coefficients $\mathbf{c}^{k}$ and $\mathbf{d}^{k}$ are explicitly given by $2k$ algebraic equations whose right-hand side depends continuously on $\mathbf{a}^{k}$ and $\mathbf{b}^{k}$. This allows us to replace $\mathbf{c}^{k}$ and $\mathbf{d}^{k}$ in the right-hand side of the aforementioned ODE system to obtain a closed ODE system for the vector-valued function $(\mathbf{a}^{k},\mathbf{b}^{k})^{\top}$ whose right-hand side depends continuously on $(\mathbf{a}^{k},\mathbf{b}^{k})^{\top}$. Consequently, the Cauchy–Peano theorem implies the existence of at least one local-in-time solution
@@
to the corresponding initial value problem. Here, we take $T_{k}^{*}>0$ as large as possible meaning that $0,T_{k}^{*})\cap[0,T]$ is the right-maximal time interval of this solution. We can now reconstruct
@@
${\boldsymbol{v}}_{k}\in C^{0}([0,T_{k}^{*});{\bf V}_{\sigma,{\bf n}}),\quad(\varphi_{k},\psi_{k})\in C^{1}([0,T_{k}^{*});{\cal V}),\quad\text{and}\quad(\mu_{k},\theta_{k})\in C^{0}([0,T_{k}^{*});{\cal V}),$
which fulfills the discretized weak formulation (3.8) on the time interval $0,T_{k}^{*})$.
@@ 3.1.2 Uniform estimates
First estimate. We first test (3.8a) by ${\boldsymbol{v}}_{k}$, (3.8b) by $\mu_{k}$, (3.8c) by $\theta_{k}$, and (3.8) by $(\partial_{t}\varphi_{k},\partial_{t}\psi_{k})$. We then add these equations and integrate the resulting equation with respect to time from 0 to an arbitrary $t\in[0,T_{k}]$. We obtain
@@
$\sqrt{2\nu_{1}}\left\|\mathbf{D}{\boldsymbol{v}}_{k}\right\|_{L^{2}(0,T_{k};{\mathbb{H}})}+\sqrt{\gamma_{1}}\left\|{\boldsymbol{v}}_{k}\right\|_{L^{2}(0,T_{k};\mathbf{H}_{\Gamma})}+\left\|\nabla\varphi_{k}\right\|_{L^{\infty}(0,T_{k};{\bf{H}})}+\left\|\nabla_{\Gamma}\psi_{k}\right\|_{L^{\infty}(0,T_{k};{\bf{H}}_{\Gamma})}$
@@
Invoking Korn’s inequality (see Lemma 2.1), we directly infer
@@
Hence, in view of (3.11) and (3.1.2), we use the Poincaré–Wirtinger inequality in $\Omega$ and Poincaré’s inequality on $\Gamma$ (see Lemma 2.2) to conclude
@@
$\quad\leq\left(\left\|\varphi_{k}\right\|_{L^{\infty}(0,T_{k};L^{6}(\Omega))}\left\|{\boldsymbol{v}}_{k}\right\|_{L^{2}(0,T_{k};{{\bf{L}}^{3}(\Omega)})}+M_{2}\left\|\nabla\mu_{k}\right\|_{L^{2}(0,T_{k};{\bf{H}})}\right)\left\|{\overline{\zeta}}\right\|_{L^{2}(0,T_{k};V)}$
@@
Third estimate. Next, we want to derive uniform bounds on $\mu_{k}$ in $L^{4}(0,T_{k};H)\cap L^{2}(0,T_{k};V)$ and on $\theta_{k}$ in $L^{4}(0,T_{k};{H}_{\Gamma})\cap L^{2}(0,T_{k};{V}_{\Gamma})$. Therefore, we choose arbitrary functions $\eta\in L^{1}(0,T_{k};V)$ and $\eta_{\Gamma}\in L^{1}(0,T_{k};{V}_{\Gamma})$ and we set ${\overline{\eta}}:=\mathbb{P}_{V_{k}}(\eta)$ and ${\overline{\eta_{\Gamma}}}:=\mathbb{P}_{V_{\Gamma,k}}(\eta_{\Gamma})$. Testing (3.8) by $({\overline{\eta}},{\overline{{\eta_{\Gamma}}}})$, recalling the growth conditions from (R1) as well as the uniform bounds (3.1.2) and (3.14), we use Hölder’s inequality and Sobolev’s embedding theorem to derive the estimate
$\left|\int_{0}^{T_{k}}\mathopen{\langle}(\mu_{k},\theta_{k}),(\eta,{\eta_{\Gamma}})\mathclose{\rangle}_{{\cal V}}\right|=\left|\int_{0}^{T_{k}}\mathopen{\langle}(\mu_{k},\theta_{k}),({\overline{\eta}},{\overline{{\eta_{\Gamma}}}})\mathclose{\rangle}_{{\cal V}}\right|$
$\quad\leq\int_{0}^{T_{k}}\Big{[}\left\|\nabla\varphi_{k}\right\|_{{\bf{H}}}\left\|\nabla{\overline{\eta}}\right\|_{{\bf{H}}}+\left\|F^{\prime}(\varphi_{k})\right\|_{L^{6/5}(\Omega)}\left\|{\overline{\eta}}\right\|_{L^{6}(\Omega)}+\left\|\nabla_{\Gamma}\psi_{k}\right\|_{{\bf{H}}_{\Gamma}}\left\|\nabla_{\Gamma}{\overline{{\eta_{\Gamma}}}}\right\|_{{\bf{H}}_{\Gamma}}$
@@
in $0,T_{k}]$. Taking the supremum over all $(\eta,{\eta_{\Gamma}})\in L^{1}(0,T_{k};{\cal V})$ with $\left\|(\eta,{\eta_{\Gamma}})\right\|_{L^{1}(0,T_{k};{\cal V})}\leq 1$, and using (3.15), we infer